a)  An electron is accelerated eastward at 1.8x10^9 m/s^2 by an electric field. Determine the magnitude and direction of the electric field.

b) Calculate the electric field strength 20 cm from a 4nC charge.


1
Expert's answer
2021-10-02T14:55:59-0400

a) The field is


"E=\frac Fq=\frac{ma}{q}=0.01\text{ N/C}."

b) The field is


"E=\frac{kq}{r^2}=899\text{ N/C}."